An alpha particle having a speed of 1.0X6^6ms^-1 collides with the stationary proton which gain an initial speed of 1.60X10^6ms^-1 in the direction of an alpha. 1. What is the speed of 2particle immediately after the collission? 2. How much kinetic energy is gain by the proton after the collission?
1
Expert's answer
2020-02-25T10:21:58-0500

1) From the conservation of momentum:

"Mv=mu-Mv'"

"v'=\frac{m}{M}u-v"

"v'=\\frac{1.67\\cdot 10^{-27}}{\u200e6.64\\cdot 10^{-27}}1.6\\cdot 10^6-1.0\\cdot 6^6"

"v'=3.56\cdot 10^5\frac{m}{s}"

2)


"K_p=0.5mu^2"

"K_p=0.5(1.67\cdot 10^{-27})(1.6\cdot 10^6)^2=2.14\cdot 10^{-15}\ J"